### v1

# SIDIS at small at next-to-leading order: gluon contribution

Abstract
We calculate the contribution of gluons to single inclusive hadron production at next-to-leading order (NLO) accuracy in Deep Inelastic Scattering (DIS) at small using the Color Glass Condensate formalism. It is shown that the only divergence present is the standard collinear divergence which is absorbed into evolution of gluon-hadron fragmentation function. Our calculations are performed at finite and we provide general finite expressions for the structure of Wilson lines appearing in inclusive dihadron and single hadron production cross sections. We also comment on how one can obtain rapidity distribution of hadron multiplicities from our results.
I Introduction
Deep Inelastic Scattering (DIS) is the cleanest environment in which to probe internal partonic structure of hadrons and nuclei Aschenauer et al. (2016); Accardi et al. (2016). Due to the experimentally observed fast rise of gluons at small Bjorken one expects that a high energy hadron or nucleus is a maximum occupancy state of predominantly gluons Gribov et al. (1983); Mueller and Qiu (1986). This high density gluonic state is referred to as the Color Glass Condensate (CGC) McLerran and Venugopalan (1994a, b); Jalilian-Marian et al. (1997a) and is expected to give the dominant contribution to scattering processes at high energy. While there have been tantalizing hints of this saturated state at HERA, RHIC and the LHC Jalilian-Marian and Kovchegov (2004); Dumitru et al. (2006); Jalilian-Marian (2006); Marquet (2007); Albacete and Marquet (2010); Stasto et al. (2012); Lappi and Mantysaari (2013); Jalilian-Marian and Rezaeian (2012a, b); Zheng et al. (2014); Stasto et al. (2018); Albacete et al. (2019); Mäntysaari et al. (2020); Hatta et al. (2021a); Jia et al. (2020); Gelis and Jalilian-Marian (2002); Dominguez et al. (2011); Metz and Zhou (2011); Dominguez et al. (2012); Iancu and Laidet (2013); Altinoluk et al. (2016a); Hatta et al. (2016); Dumitru et al. (2015); Kotko et al. (2015); Marquet et al. (2016); van Hameren et al. (2016); Marquet et al. (2018); Dumitru et al. (2019); Dumitru and Jalilian-Marian (2002a, b); Mäntysaari et al. (2019); Salazar and Schenke (2019); Boussarie et al. (2021); Ayala et al. (1996); Jalilian-Marian (2004); Kotko et al. (2017); Hagiwara et al. (2017); Henley and Jalilian-Marian (2006); Klein and Mäntysaari (2019); Hatta et al. (2021b); Kolbé et al. (2021); Gelis and Jalilian-Marian (2002); Jalilian-Marian (2005); Altinoluk et al. (2019); Boussarie et al. (2019, 2016, 2014); Dumitru and Jalilian-Marian (2010) its presence is far from being firmly established (see Iancu and Venugopalan (2003); Jalilian-Marian and Kovchegov (2006); Weigert (2005); Gelis et al. (2010); Morreale and Salazar (2021) for reviews). The proposed Electron Ion Collider (EIC) will enable us to look for gluon saturation in various channels with unprecedented precision and accuracy. Among the most promising processes in which gluon saturation is expected to play a dominant role are single and double inclusive jet and hadron production. Measurements of transverse momenta, angular correlations and rapidity dependence of these processes will shed light on dynamics of gluons saturation and its energy dependence and enable a detailed comparison with theoretical expectations based on CGC formalism Balitsky (1996); Kovchegov (2000); Jalilian-Marian et al. (1997b, 1998a, 1998b, 1999); Kovner and Milhano (2000); Kovner et al. (2000); Iancu et al. (2001); Ferreiro et al. (2002). To improve the theoretical accuracy of CGC-based predictions there has been significant progress made in calculating next to leading order (NLO) and beyond eikonal corrections to observable which are sensitive to gluon saturation Fadin and Lipatov (1998); Chirilli et al. (2012a, b); Balitsky and Chirilli (2013a, b); Grabovsky (2013); Caron-Huot (2015); Kovner et al. (2014); Lublinsky and Mulian (2017); Caron-Huot and Herranen (2018); Boussarie et al. (2018); Beuf et al. (2022a, b, 2021); Mäntysaari and Penttala (2021); Mäntysaari and Penttala (2022a, b); Lappi et al. (2022); Ayala et al. (2016, 2017, 2014); Iancu and Mulian (2021); Roy and Venugopalan (2020); Hatta et al. (2022); Iancu et al. (2022); Taels et al. (2022); Caucal et al. (2021); Bergabo and Jalilian-Marian (2022a, b); Kovner and Wiedemann (2001); Kovchegov et al. (2017a); Cougoulic and Kovchegov (2019); Kovchegov and Sievert (2019); Kovchegov et al. (2017b, c, d); Kovchegov et al. (2016); Agostini et al. (2019a, b); Altinoluk and Dumitru (2016); Altinoluk et al. (2016b); Altinoluk et al. (2014); Jalilian-Marian (2021, 2020, 2019, 2017); Hentschinski et al. (2018, 2016); Gituliar et al. (2016); Balitsky and Tarasov (2016, 2015).
In Bergabo and Jalilian-Marian (2023a) we calculated the next-to-leading order corrections to single inclusive hadron production in DIS (SIDIS) at small when either quark or antiquark in the final state hadronizes. Here we continue our work toward a complete NLO calculation of SIDIS by considering hadronization of the produced gluon 111While this manuscript was being prepared for publication we were informed of a similar work Caucal et al. (2024) which however focuses on single inclusive jet production in DIS. We thank F. Salazar for bringing their work to our attention.. This channel is not present in a Leading Order calculations of SIDIS and appears only as part of the NLO corrections so that it is important to understand its role and contribution. Furthermore we include the full dependence rather than making the large approximation.
In the small limit of DIS the virtual photon (transverse or longitudinal) splits into a quark antiquark pair (a dipole), which then multiply scatters from the target hadron or nucleus. To leading order (LO) accuracy the double inclusive production cross section can be written as
where the first and second terms in the curly bracket above correspond to the contribution of the longitudinal and transverse polarizations of the virtual photon. The virtual photon has momentum with and we have set the transverse momentum of the photon to zero without any loss of generality. Furthermore () is the momentum of the outgoing quark (antiquark) while () is its longitudinal momentum fraction relative to the photon. Also, () is the transverse coordinate of the quark (antiquark), and primed coordinates are used in the conjugate amplitude. Quark and antiquark rapidities and are related to their momentum fractions and via .
We further define and use the notation
The production cross section is a convolution of the probability for a photon to split into a quark at transverse position and an antiquark at position represented by the Bessel functions, with multiple scatterings of the quark antiquark pair on the target encoded in the dipoles and quadrupoles . To calculate the single inclusive production cross section we need to integrate over one of the final state partons which we choose to be antiquark so that quark is produced. The case when antiquark is produced is identical so that our results can be multiplied by a factor of to include antiquark production. Integrating over the produced antiquark phase space sets and gives
II Single inclusive gluon production
To calculate contribution of gluons to SIDIS we start with the next-to-leading (NLO) corrections to dihadron production in Bergabo and Jalilian-Marian (2022b). In principle there are both real and virtual contributions, however only real diagrams containing a gluon in the final state contribute so therefore we will focus on the real corrections here. The diagrams corresponding to real corrections are shown in Fig. 1,
The full NLO expressions for dihadron production are derived in Bergabo and Jalilian-Marian (2022b), in order to calculate contribution of gluons one needs to "undo" integration of the final state gluon phase space and then integrate out the final state quark and antiquark so that only the produced gluon is left in the final state. As the original expressions in Bergabo and Jalilian-Marian (2022b) are very long and not illuminating we will just show the results here,
where and are the transverse momentum and rapidity of the produced gluon with and is the fraction of the photon momentum carried by the radiated gluon . We also use the following notation for coordinate dependence of some of the Bessel functions
III Divergences
Divergences can arise due to quantum corrections to tree level cross sections. Possible divergences include Ultraviolet (UV) divergences which happen when loop -momenta go to infinity , soft divergences which arise when loop -momenta go to zero (for massless theories), collinear divergences which arise when the angle between a radiated parton and its parent (after radiation) go to zero (), and rapidity divergences when the radiated parton carries very small fraction of its parents momentum ().
It is straightforward to check that there are no UV or soft divergences in eqs. (4-13). UV divergences may occur when transverse separations (in coordinate space) go to zero, however in this limit the radiation kernels are UV finite as can be seen for example in
as , . Possible soft divergences (when transverse coordinates go to infinity in coordinate space) vanish due to behavior of Bessel functions at large values of their arguments or by the rapidly oscillating phase factors. Furthermore as the gluon is produced at finite momentum there can not be rapidity divergences. However there is a collinear divergence present in some of the terms, namely when the multiple scatterings on the target happen before radiation of the gluon (eqs. 4, 5). This collinear divergence appears as a pole in
when integrating over the quark transverse momentum (and similarly for antiquark). In principle one absorbs this collinear divergence into bare fragmentation function which makes it scale dependent. The was discussed in great detail in Bergabo and Jalilian-Marian (2022b) so here we will be brief; we define a bare parton-hadron fragmentation function and convolute the parton production cross section with this fragmentation function to describe hadronization of the parton. Next-to-leading order corrections lead to renormalization of this bare fragmentation function and make it scale dependent . Gluon contribution appears first at the NLO level so here we show it in some detail. To be specific we focus on eq. (4); we first shift and then to get
To go from the partonic to hadronic production cross section we convolute the parton level cross section with the bare gluon-hadron fragmentation function
where are now the produced hadron momenta and . We further define a new variable in terms of which the hadron production cross section can be written as
note that we now have and that . The integration over is divergent and must be regulated. Here we use dimensional regularization which gives
Further defining and then relabeling as and multiplying by to take into account contribution of antiquark radiation given by eq. (5) leads to
where we now have and the quark-gluon splitting function is defined as
This is our result for eqs. (4) and (5) which are the only terms containing divergences. At this point one can add the contribution of the LO term as well as the case of quark hadronizing already considered in Bergabo and Jalilian-Marian (2023a), the full result for terms involving collinear divergences is
Depending on the renormalization scheme (MS, ) it is standard to subtract the divergence in addition to some/no finite terms () or to cancel them using counter terms and to define the renormalized parton-hadron fragmentation function in terms of which the single inclusive hadron production cross section can be written as
where represents the finite terms given in eqs. (6-13) for gluons (and analogous but different terms for quarks Bergabo and Jalilian-Marian (2023a)). These finite contributions will be convoluted with a bare gluon-hadron (quark-hadron) fragmentation function. It is worth mentioning that if one further integrates over the produced gluon transverse momentum one can obtain the rapidity dependence of parton multiplicity distribution Caucal et al. (2024). This would then require a mechanism besides fragmentation functions to describe hadronization of the produced gluons. This may require some care specially for the collinear divergent terms in eqs. (4-5). This is however beyond the score of this work and will be investigated later.
In summary we have calculated contribution of gluons to single inclusive hadron production in Deep Inelastic Scattering at small and at finite for longitudinal photon exchange. These contributions are part of the next-to-leading order corrections to SIDIS. Some of these contributions exhibit collinear divergences which are then absorbed into Leading Log evolution of fragmentations functions. The rest are finite and constitute part of NLO contributions to SIDIS. The case of transverse photon exchange is conceptually similar but algebraically more tedious and will be reported elsewhere.
Acknowledgements:
We gratefully acknowledge support from the DOE Office of Nuclear Physics through Grant No. DE-SC0002307. We would like to thank T. Altinoluk, N. Armesto, G. Beuf, E. Iancu, Y. Kovchegov, T. Lappi, C. Marquet, F. Salazar and B. Xiao for helpful discussions. JJM would like to thank University of Santiago de Compostela, Instituto Galego de Física de Altas Enerxías (IGFAE), Jyvaskyla University, Ecole Polytechnique CPHT, Universite Paris-Saclay, IPHT and CERN for hospitality and financial support.
Appendix A Structure of finite correlations
Here we have kept the suppressed parts of NLO corrections in this work while in Bergabo and Jalilian-Marian (2022b, 2023b, 2023a) we made the large approximation. Rather than repeating the full results derived in those papers keeping finite we show the finite generalization of those expressions for the case of dihadron production. To get the finite results for the case of quark hadronizing as done in Bergabo and Jalilian-Marian (2023a) one would integrate over the final state phase space for antiquark and gluon which would then set some of these coordinates equal. Labeling of terms corresponds to the diagrams involved as before (virtual diagrams are shown in Bergabo and Jalilian-Marian (2022b, 2023b, 2023a)). We also factor out an overall factor of since the LO result also contains an explicit . These expressions then can be used to replace the large expressions in Bergabo and Jalilian-Marian (2022b, 2023b, 2023a) easily.